(i) Using 5 points x=0, 1,2,3,4 complete the table

 x     0        1       2        3        4 
f(x) 

(ii) Find "the f\u2019(0.5)."

(iii) Find the exact value if

"f\u2019(x)=(-a)\/100 e^(-ax\/100) cos\u2061\u30163ax-3a\u3017 e^(-ax\/100) sin\u20613ax"


1
Expert's answer
2021-01-12T14:43:45-0500
"Solution"

"f'(x)=\\frac{\u2212a}{100}e^{\\frac{\u2212ax}{100}}cos\u2061\u30163ax\u22123a\u3017e^{\\frac{\u2212ax}{100}}sin\u20613ax\\\\\n\\implies \\frac{d}{dx}(e^\\frac{-ax}{100}cos 3ax)"

Integrate on both sides


"f(x)=e^\frac{-ax}{100}cos\ 3ax"

Now, x=0


"f(x)=e^{-0}\ cos\ 0^0=1\\
x=1\\
f(1)=e^\frac{a}{100}\ cos\ 3a"



"\begin{matrix}
x & 0 & 1 & 2 & 3 & 4 \\ 
f(x) & 1 & e^{-\frac{a}{100}}cos\ 3a & e^{-\frac{a}{50}}cos\ 6a & e^{-\frac{3a}{100}}cos\ 9a & e^{-\frac{a}{25}}cos\ 12a
\end{matrix}"



"f'(0.5)=\frac{-a}{100}e^\frac{-a}{200}cos\ \frac{3a}{2}-3ae^\frac{-a}{200}sin\ \frac{3a}{2}\\
-ae^\frac{-a}{200}(\frac{cos \frac{3a}{2}}{100}+3\ sin\ \frac{3a}{2})"